For adiabatic ideal gas near room temperature, what fraction of heat supplied is available for external work. If the gas is expanded at a constant pressure? At a constant temperature?

1
Expert's answer
2022-01-28T12:25:25-0500

In the process of expansion at constant pressure p, assuming that the volume increases from "V_1" to "V_2" and temperature change from "T_1" to "T_2" , we have

"PV_1=nRT_1"

"PV_2=nRT_2"


In this process, the work done by the system on the outside world is "W = p(V_2-V_1)= nR\u0394T" and increase in internal energy of the system is

"\u0394U = C_v\u0394T"


"\frac{W}{Q}=\frac{W}{\Delta U + W}=\frac{nR}{C_v +nR}=\frac{2}{7}"